OpenStudy (anonymous):

find all zeros: f(x)=x^4-5x^3+10x^2-3x+3

OpenStudy (inkyvoyd):

You'll want to factor this.

OpenStudy (inkyvoyd):

Do you know synthetic division and the rational roots theorem?

OpenStudy (inkyvoyd):

@Anhvypham

OpenStudy (dumbcow):

if it factors it will be in this form: \[(x^2 + ax+3)(x^2 +bx +1)\] setting up these equations: \[a +b = -5\] \[ab + 4 = 10\] \[a + 3b = -3\]

OpenStudy (inkyvoyd):

huh. Would never have thought to approach the problem that way @dumbcow

OpenStudy (anonymous):

the answers include imaginary numbers, so that's what is throwing me offf

OpenStudy (inkyvoyd):

your equations don't seem to be consistent @dumbcow

OpenStudy (dumbcow):

@inkyvoyd , its an approach which makes it easier to get exact values it doesn't work though, function does not factor use rational root thm and synthetic division, though actually there are no real soltions

OpenStudy (inkyvoyd):

@Anhvypham where did you get this problem from? none of the roots are rational

OpenStudy (inkyvoyd):

they are also all complex.

OpenStudy (anonymous):

its in my math review packet

OpenStudy (inkyvoyd):

Well, sorry to say this, but you either made a typo or your teacher did.

OpenStudy (dumbcow):

are you given all 4 complex answers ? what class are you in?

OpenStudy (inkyvoyd):

you sure you don't mean -x^4-5x^3+10x^2-3x+3

OpenStudy (anonymous):

i have the answers. they are all complex. i'm in precalc honors

OpenStudy (inkyvoyd):

well, what are your answers?

OpenStudy (inkyvoyd):

you can reconstruct them that way

OpenStudy (dumbcow):

you could say: \[x = a \pm bi, c \pm di\] plug them in and get 4 equations and try to solve for a,b,c,d its really messy though

OpenStudy (inkyvoyd):

I was thinking for complex roots a, b, c, and d he could just take 0=(x-a)(x-b)(x-c)(x-d) on a CAS system to find out the actual correct equation, then resolve the actual correct equation

OpenStudy (inkyvoyd):

if you give me the roots I can give you the equation that it expands to.

OpenStudy (inkyvoyd):

though it kinda loses the meaning with the answers since you know what to do

OpenStudy (anonymous):

roots: 3i, -3i, \[\frac{ (3i \sqrt{2}) }{ 2 }\], \[-\frac{ (3i \sqrt{2}) }{ 2 }\]

OpenStudy (dumbcow):

whoa ok you have a typo in your function

OpenStudy (inkyvoyd):

well that doesn't seem to check out at all. Your roots are all imaginary?

OpenStudy (anonymous):

yea

OpenStudy (inkyvoyd):

put that in wolfram alpha and you get x^4+(27 x^2)/2+81/2.... which has nothing to do with your original equation. lol.

OpenStudy (anonymous):

this problem makes no sense. it probably was a typo. thanks for your help though guys

OpenStudy (inkyvoyd):

yeah I was hoping that through the roots you could reconstruct what the problem was supposed to be, but I guess not lol